Hello!
The number of a certain species of fish is modeled by the function below where t is measured in years and n(t) is measured in millions.
n(t) = 11e^0.014t
After how many years will the number of fish reach 30 million? (Give your answer correct to one decimal place.?
thanks for your help!

Answer by RAY100(1637)

30,000,000 = 11 e^ (.014t)
.
divide by 11
.
2,727.272.7 = e^ (.014T)
.
take ln of both sides, remember ln e^a = a
.
14.8188 = .014 t
.
divide by .014
.
1058.4866 = t
.
checking
.
11 * e ^( .014 * 1058.4866) = 30,000,000,,,,,ok